the perimeter of a rectangle is 64cm.the length is 4 cm. less than twice the width.find the length and the width.
Answer by ankor@dixie-net.com(22740)

:
the perimeter of a rectangle is 64 cm.
2L + 2W = 64
:
the length is 4 cm. less than twice the width.
L = (2W-4)
:
find the length and the width.
:
Substitute (2W-4) for L in the perimeter equation:
2(2W-4) + 2W = 64
:
Simplify, divide equation by 2:
(2W-4) + W = 32
2W - 4 + W = 32
2W + W = 32 + 4
3W = 36
W = 36/3
W = 12 cm
:
Find the Length using L = 2W-4:
L = 2(12) - 4
L = 20 cm
:
:
Check: 2(20) + 2(12) = 64